OpenStudy (babyslapmafro):

Please show me where I went wrong with solving the following problem (click to see calculations). Find the volume of the solid that results when the region enclosed by x=y^2 and x=y is revolved about the line y=-1

OpenStudy (babyslapmafro):

\[\pi \int\limits_{0}^{1}[(\sqrt{x}-(-1))^2-(x-(-1))^2]dx\] \[=\pi \int\limits_{0}^{1}[x-x^2]dx=\pi[\frac{ x^2 }{ 2 }-\frac{ x^3 }{ 3 }]^{1}_{0}\]

OpenStudy (babyslapmafro):

\[=\frac{ \pi }{ 6 }\]

OpenStudy (babyslapmafro):

the answer in the book is pi/2

OpenStudy (babyslapmafro):

@zepdrix

OpenStudy (precal):

you can still edit your question

OpenStudy (babyslapmafro):

thanks

OpenStudy (precal):

bare with me just checking your math from your set up